# AI forecasting of higher-order wave modes of spinning binary black hole mergers

### v1

@@ 2.2.3 Decoder
The decoder receives both the output from the encoder as well as the decoder input, which has been embedded in the same way as the encoder input. First, a masked multihead self-attention mechanism is applied to the decoder input only. This mechanism prevents each position in the decoder input from attending to subsequent positions in the sequence, preserving causality and enabling autoregressive predictions during inference. The operation is defined as:
@@
$M$ is a mask matrix that applies negative infinity to positions not to be attended to, ensuring that the softmax operation assigns them zero weight. As before, $Q^{j}$, $K^{j}$, and $V^{j}$ are the query, key, and value matrices derived from linear transformations on the decoder’s input. The dimensionalities of these matrices are the same as those of the encoder self-attention matrices with the exception of the sequence length $n$, which varies for encoder inputs and decoder inputs. Outputs from each attention head are again concatenated and linearly transformed. These transformations leave the shape of the decoder input unchanged.
Following the self-attention layer, the cross-attention layer allows the decoder to focus on relevant positions in the encoder output sequence. The queries $Q$ stem from the previous decoder layer (denoted as $Y\in\mathbb{R}^{n_{dec}\times d_{embed}}$), while the keys $K$ and values $V$ are derived from the encoder output $X$:
@@
This operation preserves the shape of the queries, in this case the decoder input, ensuring the sequence length of the decoder output matches that of the decoder input. Again, $QK^{T}$ serves as a measure of similarity between keys and queries. The array returned by the softmax activation function is of shape $\mathbb{R}^{n_{dec}\times n_{enc}}$ and serves as an array of weights for the value array $V$. Figure 8 presents a visualization of these weights for a choice of input waveform and attention head. The subsequent processing after application of the attention mechanisms mirrors that of the encoder module: each position’s output is fed to a feed-forward network, residual connections and layer norms are applied after both attention layers as well as the feed forward network to stabilize training. Finally, a 1D convolution is applied to the decoder output to produce predictions for the two time-series $(h_{+},h_{\times})$ polarizations, respectively.
@@ 2.2.4 $h_{\times}$ mask
Empirically, we find that separating waveforms with null cross-polarization leads to slightly higher reconstruction accuracy. To facilitate this, we apply a mask to identify and filter waveforms with inclination angle $\theta=\pi/2$, as these exhibit null cross-polarization. This mask allows us to modify the final convolutional layer and feed-forward layers, which are specifically separated for these waveforms to better capture their unique characteristics. The remaining transformations in the model are consistent across all waveforms, ensuring uniform processing while accommodating the distinct features of the null cross-polarization cases.
3 Training and inference
@@
For our transformer model, we used an embedding dimension of 160 (this includes both the contributions from the real and the imaginary part), a feed-forward dimension of 80, and 10 attention heads. This model was trained on the NCSA Delta system, using 4 quad NVIDIA A100 GPU nodes. We employed a batch size of 16 and trained the model on 14,440,761 waveforms using the Adam optimizer with mean squared error (MSE) loss metric. The initial learning rate was set to 0.001, with decay applied during training. Training was distributed across all GPUs using PyTorch’s DistributedDataParallel (DDP) framework. The model was trained for 30 epochs and reached convergence in around 15 hours.
@@
Inference on 840,000 test waveforms was completed within 7 hours using 1 node on the DeltaAI system, i.e., 4 NVIDIA H100 GPUs. During training, the model processes entire sequences at once. Inference, in contrast, involves sequential autoregressive generation, where each output is used to predict the next timestep. The functional differences between training and inference are illustrated in Figure 5.
4 Results and Discussion
We evaluated the performance of the transformer model on a test set of 840,000 waveforms that covers the signal manifold under consideration. We used the overlap, ${\cal{O}}$, as the figure of merit to quantify the accuracy of forecasted waveforms. This metric is equivalent to the cosine similarity between the predicted and target waveforms. Specifically:
@@ 4 Results and Discussion
${\cal{O}}(h_{\mathrm{true}},h_{\mathrm{pred}})$ $=\underset{t_{c}}{\mathrm{max}}\left(\frac{\langle h_{\mathrm{true}},h_{\mathrm{pred}}[t_{c}\,]\rangle}{\sqrt{\langle h_{\mathrm{true}},h_{\mathrm{true}}\rangle\langle h_{\mathrm{pred}}[t_{c}\,],h_{\mathrm{pred}}[t_{c}\,]\rangle}}\right)\,,$ (10)
where $\langle h_{\mathrm{true}},h_{\mathrm{pred}}\rangle$ denotes the inner product between the complex-valued waveforms $h_{\mathrm{true}}$ and $h_{\mathrm{pred}}$, and $h_{\mathrm{pred}}[t_{c}]$ indicates that the waveform has been time-shifted. A selection of waveform comparisons is depicted in Figure  6, demonstrating the agreement between predicted and target waveforms for various parameter values. The figure includes two randomly selected waveforms, an example of a low overlap waveform, and a waveform with $\theta=\pi/2$, where the cross polarization $h_{\times}=0$.
@@
To provide a more detailed assessment of the model’s performance, the top panel in Figure 7 presents a histogram of the overlap values across the test set. Notably, none of the overlap values fall below 0.85, only $0.003\%$ of test set waveforms fall below 0.9, while $7.15\%$ fall below 0.99, indicating that the model achieves high accuracy in most cases. The mean and median overlaps are 0.996 and 0.997, respectively.
The bottom panels in Figure 7 display overlaps as a function of merger parameters, specifically mass ratio $q$, inclination angle $\theta$, and effective spin $s^{z}_{eff}$, defined as [43]:
$s_{\textrm{eff}}^{z}=\frac{qs^{z}_{1}+s^{z}_{2}}{1+q}\,.$ (11)
Areas of comparatively poor reconstruction accuracy are those at the edge of the signal manifold, i.e., both low mass ratio, $q$, and low inclination angle, $\theta$, as well as those of both high mass ratio and inclination angle around $\pi/2$, as observed in the bottom left panel. Additionally, we observe that low spin values of the heavier object, $s^{z}_{1}$, generally result in slightly lower reconstruction accuracy, with these regions visible as distinctive green bands in the bottom right panel.
@@
Figure 7: Top panel: Histogram of overlaps. Bottom panels: Overlap $\cal{O}$ as a function of merger parameters. Point opacity is inversely proportional to overlap value, with overlaps close to 1 rendered fully transparent, to highlight poor overlaps that might otherwise be obscured. A slight random jitter has been added for the same reason.

### v2

@@ 2.2.3 Decoder
The decoder receives both the encoder output and the decoder input, which is embedded using the same method as the encoder input. First, a masked multi-head self-attention mechanism is applied to the decoder input only. This mechanism prevents each position in the decoder input from attending to subsequent positions in the sequence, preserving causality and enabling autoregressive predictions during inference. The operation is defined as:
@@
where $M$ is a mask matrix that applies negative infinity to positions not to be attended to, ensuring that the softmax operation assigns them zero weight. As before, $Q^{j}$, $K^{j}$, and $V^{j}$ are the query, key, and value matrices derived from linear transformations on the decoder’s input. The dimensionalities of these matrices are the same as those of the encoder self-attention matrices, with the exception of the sequence length $n$, which varies for encoder inputs and decoder inputs. Outputs from each attention head are again concatenated and linearly transformed. These transformations leave the shape of the decoder input unchanged.
Following the self-attention layer, the cross-attention layer allows the decoder to focus on relevant positions in the encoder output sequence. The queries $Q$ stem from the previous decoder layer (denoted as $Y\in\mathbb{R}^{n_{\text{dec}}\times d_{\text{embed}}}$), while the keys $K$ and values $V$ are derived from the encoder output $X$:
@@
This operation preserves the shape of the queries, in this case the decoder input, ensuring the sequence length of the decoder output matches that of the decoder input. $QK^{T}$ serves as a measure of similarity between keys and queries. The array returned by the softmax activation function is of shape $\mathbb{R}^{n_{\text{dec}}\times n_{\text{enc}}}$ and serves as an array of weights for the value array $V$. Figure 12 presents an instructive visualization of these weights for a choice of input waveform and attention head.
The subsequent processing after the application of the attention mechanisms mirrors that of the encoder module: each position’s output is fed to a feed-forward network, residual connections, and layer norms are applied after each attention layer and the feed-forward network to stabilize training. Finally, a 1D convolution is applied to the decoder output to produce predictions for the two time-series $(h_{+},h_{\times})$ polarizations, respectively.
@@ 2.2.4 $h_{\times}$ Mask
Empirically, we find that separating waveforms with null cross-polarization leads to slightly higher reconstruction accuracy. To facilitate this, we apply a mask to identify and filter waveforms with inclination angle $\theta=\pi/2$, as these exhibit null cross-polarization. This mask allows us to modify the final convolutional layer and feed-forward layers, which are implemented separately for these waveforms to better capture their distinct waveform structure. The remaining transformations in the model are consistent across all waveforms, ensuring uniform processing while accommodating the distinct features of the null cross-polarization cases. The quantitative impact of this mask is outlined in section 3.5.
2.3 Training and Inference
@@
For our transformer model, we used an embedding dimension of 160 (this includes both the contributions from the real and the imaginary part), a feed-forward dimension of 80, and 10 attention heads. The model is trained on a dataset of 14,440,761 waveforms, each consisting of 5,130 total timesteps, with only 362,722 trainable parameters. This results in a high data-to-parameter ratio, which may improve generalization capabilities and help prevent overfitting, as the model is exposed to orders of magnitude more training samples relative to its size.
@@
This model was trained on the NCSA Delta system, using 4 quad NVIDIA A100 GPU nodes. We employed a batch size of 16 and trained the model using the Adam optimizer [35] with a mean squared error (MSE) loss metric. The initial learning rate was set to 0.001, with decay applied during training. Additional dropout or weight decay was not required. Training was distributed across all GPUs using PyTorch’s DistributedDataParallel (DDP) framework [36]. The model was trained for 30 epochs and reached convergence in around 15 hours.
@@
Inference on 840,000 test waveforms was completed in under 5 hours using 1 NVIDIA V100 GPU on NCSA’s HAL system. During training, the model processes entire sequences at once. Inference, by contrast, proceeds via sequential autoregressive generation, with each output timestep used to predict the next timestep, which is in turn fed back into the transformer decoder module. The functional differences between training and inference are illustrated in Figure 5. During inference, the encoder outputs remain constant for each input signal and are therefore computed only once at the beginning of the autoregressive prediction. This approach substantially reduces processing time.
2.4 Forecasting Efficiency
To evaluate the inference performance of our transformer-based sequence model for gravitational wave signals, we benchmarked it against the surrogate waveform model NRHybSur3dq8, focusing on waveform generation latency. We selected 50 representative test points within the binary black hole parameter space using a Halton sequence, which ensures uniform coverage and reduces clustering compared to purely random sampling[37]. At each point, we generated waveforms 10 times with both models and recorded the per-point average. The surrogate model was run on a single AMD Milan CPU core on the NSCA Delta cluster [^1] For the surrogate, performance remained consistent for waveform start times as early as -12000M, consistent with prior benchmarks [38]., while the transformer model was benchmarked on a single NVIDIA A100 GPU on the same cluster. On average, the surrogate generated one waveform in 151 ms (6.6 waveforms/s), and the transformer completed inference for one waveform in 18 ms (55.6 waveforms/s). [^2] Assuming ideal parallel scaling on one Delta node consisting of AMD EPYC 7763 “Milan” processors with 64 cores and 4 NVIDIA A100 GPUs, these single-instance results suggest a maximum throughput of approximately 422 waveforms per second for the surrogate using all 64 CPU cores, and  222 waveforms per second for the transformer using all 4 NVIDIA A100 GPUs.
The surrogate model reconstructs waveforms through interpolation of precomputed numerical relativity data, making it computationally lightweight and efficient within its training domain. In contrast, the transformer performs autoregressive sequence forecasting based on learned dynamics of strong-field spacetime evolution, enabling it to generalize beyond the training distribution. While this introduces greater computational complexity per forecast, it also allows the transformer to adapt to previously unseen input conditions. The transformer additionally benefits from GPU-optimized tensor operations and efficient batch processing, whereas the surrogate is more memory-bound and exhibits diminishing returns when parallelized beyond core saturation.
3 Results
We evaluated the predictive accuracy of our transformer-based sequence model on a test set comprising 840,000 waveforms spanning the full extent of the targeted signal manifold. The accuracy of generated waveforms is quantified using the overlap metric, $\mathcal{O}$, which captures the normalized inner product between predicted and true waveforms, maximized over a time shift:
@@ 3 Results
$\mathcal{O}(h_{\mathrm{true}},h_{\mathrm{pred}})=\underset{t_{c}}{\mathrm{max}}\left(\frac{\langle h_{\mathrm{true}},h_{\mathrm{pred}}[t_{c}]\rangle}{\sqrt{\langle h_{\mathrm{true}},h_{\mathrm{true}}\rangle\langle h_{\mathrm{pred}}[t_{c}],h_{\mathrm{pred}}[t_{c}]\rangle}}\right),$ (10)
[…]
3.1 Visual Inspection of Predictions
Figure 6 presents a small sample of model outputs across a range of parameter regimes, compared to the surrogate-based ground truth. The figure includes: (i) two randomly selected waveforms, (ii) one with relatively low overlap, and (iii) one at $\theta=\pi/2$, where $h_{\times}=0$. These examples provide a visual reference for the range of prediction quality observed across the dataset. For an interactive gallery of more predicted and true waveforms, visit  [29].
@@ 3.1 Visual Inspection of Predictions
3.2 Overall Performance Across the Signal Manifold
[…]
@@ 3.2 Overall Performance Across the Signal Manifold
Figure 7: Top panel: Histogram of overlaps. Bottom panels: Overlap $\cal{O}$ as a function of merger parameters. Point opacity is inversely proportional to overlap value, with overlaps close to 1 rendered fully transparent to highlight poor overlaps that might otherwise be obscured. A slight random jitter has been added for the same reason.
3.3 Performance on Numerical Relativity Waveforms
We conducted an additional benchmark using numerical relativity waveforms from the SXS catalog[39, 40, 41]. Specifically, we selected all available quasi-circular, non-precessing black hole merger simulations, and reconstructed the full strain signal (excluding memory modes) using all available spherical harmonic modes. While our model was trained only on a subset of modes (see section 2.1) and within a limited parameter space ($q\leq 8$ and $|s_{\{1,2\}}^{z}|\leq 0.8$), the NR dataset spans a broader range (including systems with $q\leq 15$ and $|s_{\{1,2\}}^{z}|\leq 1.0$). We project each of the simulations using a random inclination angle $\theta\in[0,\pi]$. We retained only those waveforms that extend to at least $-5000\textrm{M}$ before merger to match the encoder input requirements, leaving 521 waveforms.
Despite the mismatch in training domain and harmonic content, the model generalizes reasonably well: we find a median overlap of $\cal{O}$ = 0.969 and a mean overlap of $\cal{O}$ = 0.902 between the transformer predictions and the full NR waveforms. As expected, these scores are lower than the surrogate-based test set, but remain high given the broader parameter coverage.
We find that low-overlap outliers are broadly distributed across inclination angles, but are absent near face-on or face-off configurations ($\theta\approx 0\text{ or }\pi$). Restricting to systems within $\pi/4$ of those orientations raises the median overlap to 0.990; further narrowing to $\pi/8$ yields a median of 0.998 and a minimum above 0.95—demonstrating exceptional fidelity in detector-favorable regimes. Notably, at any inclination angle, lower overlaps are not concentrated in out-of-distribution regions.
For a baseline comparison to our transformer-based forecasting model, we consider another machine learning model that reconstructs gravitational waveforms directly from source parameters. While most machine learning models for gravitational waveforms report performance on surrogate data or synthetic signals, few are evaluated directly on full numerical relativity waveforms. One exception is NRSurNN3dq4 [42], a neural network surrogate trained on the $(2,2)$ mode of aligned-spin, non-precessing BBH waveforms with mass ratios up to $q=6$ and dimensionless spin components $\in[-0.99,0.99]$, using the NRSur7dq4 surrogate [43]. It was then fine-tuned and evaluated on a filtered set of 381 SXS waveforms within the same parameter range, achieving a mean overlap of 0.995 and a minimum overlap of 0.979 on a held-out NR test set. Their model differs from ours in mode content, parameter space, and modeling objective (parameter-to-waveform mapping versus time-series forecasting).
Our results suggest that the model retains considerable generalization capacity even when applied to physical systems and mode structures beyond those it was explicitly trained on.
3.4 Dependence on Physical Parameters
To examine how predictive accuracy on the surrogate-based test set varies across the parameter space, we first visualize overlap scores as a function of key physical parameters. The bottom panels of Figure 7 display two scatter plots: one mapping overlap against mass ratio and inclination angle $(q,\theta)$, and the other against mass ratio and effective spin $(q,s_{\mathrm{eff}}^{z})$. In both plots, point opacity is inversely scaled with overlap value, making regions of lower reconstruction accuracy visually prominent. These visualizations reveal several trends:
  • Mass ratio ($q$): Degradations are observed at the edges of the parameter space for $q\sim\{1,8\}$.
  • Inclination angle ($\theta$): Reduced performance is localized near $\theta\sim\{0,\pi/2\}$, corresponding to face-on and edge-on configurations, respectively.
  • Effective Spin ($s_{\mathrm{eff}}^{z}$): Defined as
    $s_{\mathrm{eff}}^{z}=\frac{qs_{1}^{z}+s_{2}^{z}}{1+q},$ (11)
    a standard parameter used in gravitational-wave population modeling [44]. Highly anti-aligned spins (i.e., lower values of $s_{1}^{z}$, the spin component of the primary black hole) are associated with a reduction in reconstruction quality, with these regions visible as distinctive green bands in the bottom right panel of Figure 7.
We provide additional results in A, where we study the behavior of the transformer by partitioning the signal manifold into fixed spin slices across the $(q,\theta)$ plane.
3.5 Importance of the $h_{\times}$ Masking Mechanism
An important component of this model is the application of a conditional masking mechanism that affects edge-on orientations where the cross polarization $h_{\times}\approx 0$. Ablation tests excluding this mechanism resulted in a decrease in average overlap to 0.994, and a drop in minimum overlap to 0.585 on the surrogate-based test set. These results underscore the importance of the $h_{\times}$ mask for maintaining robustness across observational scenarios. Excluding this conditional masking mechanism yields a $\sim 15\%$ inference speedup, offering a trade-off that can be tuned based on application requirements.
3.6 Exploratory Interpretability Studies
These analyses aim to identify which waveform features the model relies on most during prediction. While interpretability results may reflect aspects of the input encoding or architecture—such as symmetry, vanishing components, or decoder proximity—many of the observed patterns align with known physical principles in gravitational wave modeling. The following studies should thus be viewed as indicative of learned model behavior, with findings potentially offering insight into physically meaningful signal properties.
3.6.1 Obfuscation Experiments
To understand which regions of the waveform most contribute to predictive performance, we conduct a systematic input obfuscation study. In each test, we zero out a fixed segment of the encoder input—effectively masking that region—and measure the resulting degradation in predictive performance. This allows us to estimate the relative importance of each input interval. For a gallery of predicted waveforms resulting from these obfuscation experiments, see [29].
Figure 8 illustrates the masking procedure: a segment of the encoder input (blue) from $t\in[-2000\mathrm{M},-1000\mathrm{M}]$ is replaced with zeros (red), while the decoder target (orange) remains unaltered. This setup ensures that any change in output quality can be directly attributed to the missing input region.
Figure 8: Example of obfuscation: a segment of the input waveform between $t\in[-2000\mathrm{M},-1000\mathrm{M}]$ is zeroed out before inference.
[…]
To probe how this sensitivity varies with physical parameters, we calculate the Pearson correlation coefficient between the overlap drop (i.e., performance loss from masking) and each system’s intrinsic properties: mass ratio $q$, spin components $(s_{1}^{z},s_{2}^{z})$, and inclination angle $\theta$. Since inclination effects are approximately symmetric around $\theta=\pi/2$, we use the transformed variable $\lvert\theta-\pi/2\rvert$ to better reflect this structure. The resulting correlation matrix is shown in Figure 10. While correlation values provide a coarse summary of parameter importance, they can obscure nonlinear relationships. B shows binned overlap drops, revealing structure not captured by Pearson correlation alone.
Figure 9: Distribution of overlap scores $\mathcal{O}$ for different obfuscated input intervals. Each box shows the interquartile range, with the median marked by a central line. Whiskers extend to 1.5×1.5$\times$1.5 × IQR, and outliers are shown as individual points.
The correlations reveal several noteworthy trends:
  • Mass Ratio ($q$): A moderate negative correlation appears in the earliest segments, suggesting that equal-mass binaries carry more predictive utility in the early inspiral compared to asymmetric systems. While high mass-ratio systems exhibit richer harmonic content due to enhanced subdominant modes, equal-mass systems produce smoother, more symmetric signals dominated by the quadrupole mode [45]. These smoother signals may be more readily captured and generalized by the network in this regime. The features of asymmetric binaries become more prominent, and more predictive, closer to merger, as reflected in the steadily increasing match drop with $q$ in later segments.
  • Spin of the More Massive Black Hole ($s_{1}^{z}$): The correlation with overlap drop is strongest in the final input segment ($-200\textrm{M},-120\textrm{M}$), suggesting that the model relies more heavily on late-time features when $s_{1}^{z}$ is aligned with the orbital angular momentum. This is in line with previous works which show that spin effects are most prominent near merger, where relativistic velocities enhance spin–orbit and spin–spin couplings. These contribute at higher post-Newtonian orders and thus have limited influence during early inspiral. Near merger, high spins affect phase evolution and waveform amplitude, as reflected in the amplitude’s dependence on final spin and the emergence of spin-driven multipolar radiation [46, 47].
  • Spin of the Less Massive Black Hole ($s_{2}^{z}$): Correlations with $s_{2}^{z}$ are weak across all intervals, indicating a diffuse and relatively minor influence on waveform morphology. This is consistent with post-Newtonian predictions that attribute only a minor role in phase evolution to the spin of the smaller object [48].
    Figure 10: Correlation matrix showing how waveform parameters correlate with the drop in overlap when different parts of the input are obfuscated. Each row corresponds to a physical parameter: mass ratio $q$, spin components $s_{1}^{z}$ and $s_{2}^{z}$, and the transformed inclination $|\theta-\frac{\pi}{2}|$, which captures the approximately symmetric structure observed around edge-on orientations. Each column corresponds to a time window in units of $\mathrm{M}$.
  • Inclination Angle ($\theta$): The inclination angle exhibits nontrivial correlation even in early inspiral segments, suggesting that the model leverages amplitude and polarization modulations induced by viewing geometry across long timescales. Such modulations arise because inclination affects how spherical harmonic modes project onto the line of sight: face-on systems are dominated by the $(2,2)$ mode, while edge-on systems excite a broader spectrum of harmonics. These geometric effects alter signal morphology across the entire waveform, as noted in [49]. Notably, the match drop for edge-on binaries is near zero except in the final segments $(-1000\mathrm{M},-120\mathrm{M})$ (see B), likely due to the suppressed cross-polarization component at $\theta\approx\pi/2$, which reduces the effect of masking rather than indicating model disinterest. As such, low match drops near edge-on inclination should be interpreted with caution, as they may reflect vanishing waveform content rather than reduced importance.
Overall, these results offer quantitative insight into the localized importance of waveform features across time. Across all binary configurations, aligned-spin systems (i.e., systems where the primary black hole’s spin is aligned with the orbital angular momentum) and those with large $\theta\approx\pi$ are among the best resolved by the model and exhibit the highest information content in the late inspiral, as revealed by our obfuscation studies. In contrast, systems with anti-aligned primary spin or those with $\theta\approx 0$ carry less predictive information in the final input segments and correspond to regions of degraded performance. These patterns suggest a possible relationship between the informational richness of the late inspiral and the model’s reconstruction accuracy: systems where specifically the late-time waveform evolution is more informative tend to be more successfully forecasted.
3.6.2 Mode Modified Datasets
To evaluate the contribution of higher-order gravitational wave modes to model accuracy, we constructed two additional test sets with truncated mode content. The first includes only modes up to $\ell_{\max}=3$, while the second is limited to $\ell_{\max}=2$. These reduced-mode datasets are generated using the same procedure and cover the same temporal domain as the full dataset, $t\in[-5000\textrm{M},130\textrm{M}]$, but exclude subdominant modes. Unlike in the obfuscation experiments, the ground truth here is adjusted to match the ablated input. This ensures that performance differences reflect the absence of subdominant modes rather than inconsistencies between input and target.
For each dataset, we compute the overlap score $\mathcal{O}$ between model predictions and the corresponding reduced-mode ground truth. Summary statistics are presented in Table 1. As anticipated, the full-mode dataset, $\ell_{\max}=4$, yields the highest mean and median overlaps and the lowest standard deviation, whereas reducing mode content progressively decreases accuracy and increases variance.
Table 1: Overlap statistics for different mode cutoffs.
Dataset Mean $\mathcal{O}$ Median $\mathcal{O}$ Std
$\ell_{\max}=2$ 0.9687 0.9890 0.0575
$\ell_{\max}=3$ 0.9864 0.9938 0.0226
$\ell_{\max}=4$ (Full) 0.9959 0.9974 0.0052
To understand where these discrepancies originate, we analyze the performance drop relative to the full-mode dataset as a function of intrinsic binary parameters. Figure 11 displays the loss in overlap for each reduced-mode variant across mass ratio $q$, the aligned spin of the heavier black hole $s^{z}_{1}$, and inclination angle $\theta$.
Figure 11: Each point compares either ${\mathcal{O}}\left(\ell_{\max}=2\right)-{\mathcal{O}}\left(\ell_{\max}=4\right)$ or ${\mathcal{O}}\left(\ell_{\max}=3\right)-{\mathcal{O}}\left(\ell_{\max}=4\right)$ as a function of binary parameters. The inclusion of higher order modes improves forecasting accuracy for binaries with large mass ratios, low spin values of the heavier black hole, $s^{z}_{1}$, and edge-on orientations.
Our analysis reveals that higher-order modes improve accuracy in three specific regimes:
  • High Mass Ratio ($q\gtrsim 4$): The exclusion of higher modes leads to significant overlap loss, particularly when reducing to $\ell_{\max}=2$. This aligns with the increased contribution of subdominant harmonics in asymmetric mergers. As mass ratio deviates from 1, symmetry suppression of odd-$m$ and $-\ell$ modes weakens, and subdominant modes grow in significance. Mode truncation most affects high-$q$ systems, where the $(2,2)$ mode no longer dominates [45].
  • Anti-Aligned Primary Spin ($s^{z}_{1}<0$): Systems where the primary black hole’s spin is strongly anti-aligned with the orbital angular momentum show somewhat increased sensitivity to mode reduction, particularly in a small subset of cases with $s^{z}_{1}\ll 0$ when excluding $\ell>3$. This trend is modest, and further analysis indicates it does not strongly correlate with spin–spin alignment or with effective spin. The increased mode sensitivity may instead reflect a compensatory role of subdominant harmonics in systems with retrograde alignment, which are known to exhibit distinct merger dynamics and reduced inspiral duration compared to aligned configurations [50, 51].
  • Edge-On Orientations ($\theta\approx\pi/2$): Orientation effects are especially pronounced in the transition from $\ell_{\max}=4$ to $\ell_{\max}=3$, but not between $\ell_{\max}=3$ and $\ell_{\max}=2$. This suggests that modes with $3<\ell\leq 4$ encode angular structure essential for modeling polarization and amplitude variations in edge-on configurations. For such orientations, the dominant $(2,2)$ mode contributes less power along the line of sight, while subdominant modes project more strongly. As a result, neglecting higher-order modes leads to the largest mismatches in edge-on cases, where accurate waveform modeling requires their inclusion [49].
In contrast, the spin of the lighter black hole $s^{z}_{2}$ continues to play a secondary role across all mode configurations. The consistently weak dependence on $s^{z}_{2}$ suggests that the model captures its influence through more global, lower-order features.
Overall, these results confirm that including higher-order modes improves model fidelity, particularly in regimes where subdominant harmonics contribute significantly to the waveform structure. Systems where higher-order modes are essential—such as high mass-ratio, edge-on mergers—exhibit both high sensitivity to mode removal and reduced baseline performance, suggesting that dependence on subdominant mode content correlates with reconstruction difficulty, likely due to the increased structural complexity and reduced dominance of the quadrupole mode in these regimes. The model’s sensitivity to angular effects and binary asymmetries further highlights the value of full mode content for achieving robust and generalizable predictions.
3.7 Attention Plots